A garden hose has a radius of 0.0120 m, and water comes out at a speed of 4.88 m/s. How much time does it take to fill up a kids swimming pool with a volume of 3.88 m^3
1
Expert's answer
2018-05-22T10:02:08-0400
dV/dt=Av=πr^2 v
V=πr^2 vt
t=V/(πr^2 v)=3.88/(π(0.0120)^2 4.88)=1760 s.